Calculus I, Fall 2015 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Week 15

1. Section 7.3: Problems 1, 11, 15, 19, 37, 45, 57.

2. Section 7.3: Derive Table 7.8 using formula for derivatives of inverse functions.

3. Section 7.4: Problems 7, 8, 9, 10, 23(a), 24.

4. Chap 7, Additional and Advanced Exercises: problems 6.